OpenStudy (anonymous):

factor by grouping 6x^3-6x^2-x+1

OpenStudy (anonymous):

6x^3 - 6x^2 - x + 1 = 6x^2 (x -1) -1(x - 1) = (x-1) (6x^2 - 1)

OpenStudy (ash2326):

isuc did you understand?

OpenStudy (anonymous):

not really but i'm trying to

OpenStudy (anonymous):

I don't get where all the -1's are coming from

OpenStudy (ash2326):

Ok let me help you we have \[6x^3-6x^2-x+1\] the first term can be written as \(6x^3=x\times 6x^2\) let's replace that \[6x\times 6x^2-6x^2-x+1\] let's take out that 6x^2 common from first two terms \[6x^2(x-1)-x+1\] Did you understand till now?

OpenStudy (anonymous):

I don't but it's not your fault, I just don't understand where the -1 is coming from still lol sorry but I guess I'm a little slow

OpenStudy (ash2326):

Now worries I'll teach you If i Have 1 it can be written also as \(-1\times -1\) right?

OpenStudy (ash2326):

*No

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh okay

OpenStudy (ash2326):

Now we have \[6x^2(x-1)-x+1\] I can write -x-1 as -1(x-1) so \[6x^2(x-1)-(x-1)\] this "-" is actually a -1 Do you get this?

OpenStudy (anonymous):

I see now, so you just did 6x^2(x-1) because 6x^2 *x = 6x^3 and 6x^2*-1 is -6x^2

OpenStudy (ash2326):

Yeah:) now we can take x-1 out \[6x^2(x-1)-1(x-1)\] SO \[(x-1)(6x^2-1)\] do you understand now?

OpenStudy (anonymous):

man I feel dumb lol thank you for the help! I appreciate you taking the time to explain!

OpenStudy (anonymous):

not that you made me feel dumb just that I realize now how simple it is.

OpenStudy (katrinakaif):

Smart isn't something you are; It is something you become

OpenStudy (ash2326):

No worries, glad to help. @Isuckatmath1 why don't you change your user name to something like wannalearnmath:)

OpenStudy (katrinakaif):

It is better to view Math with an outlook of optimism rather than through the spectacles of pessimism

OpenStudy (anonymous):

lol I suppose I could do that if it lets me.

OpenStudy (ash2326):

make a new account :)

OpenStudy (anonymous):

thank you also katrina for the words of wisdom, and I just might do that Ash. Thanks again, I do really appreciate it

OpenStudy (katrinakaif):

Anytime @Isuckatmath1